# A Multi-Stream Fusion Approach with One-Class Learning for Audio-Visual Deepfake Detection

### v1

@@ Abstract
This paper addresses the challenge of developing a robust audio-visual deepfake detection model. In practical use cases, new generation algorithms are continually emerging, and these algorithms are not encountered during the development of detection methods. This calls for the generalization ability of the method. Additionally, to ensure the credibility of detection methods, it is beneficial for the model to interpret which cues from the video indicate it is fake. Motivated by these considerations, we then propose a multi-stream fusion approach with one-class learning as a representation-level regularization technique. We study the generalization problem of audio-visual deepfake detection by creating a new benchmark by extending and re-splitting the existing FakeAVCeleb dataset. The benchmark contains four categories of fake video(Real Audio-Fake Visual, Fake Audio-Fake Visual, Fake Audio-Real Visual, and unsynchronized video). The experimental results show that our approach improves the model’s detection of unseen attacks by an average of 7.31% across four test sets, compared to the baseline model. Additionally, our proposed framework offers interpretability, indicating which modality the model identifies as fake.
@@ I Introduction
Recent research primarily focuses on fusing features from both modalities to improve the detection performance on audio-visual deepfake datasets [9, 10, 11]. By leveraging the complementary nature of audio and visual data, these approaches effectively improve their accuracy in identifying manipulated content. However, two issues are not well-explored: First, the existing deep learning models may overfit to the specific fake generation methods present in the training data, leading to poor generalization when confronted with unseen deepfake generation algorithms in real-world scenarios. This could be attributed to the existing dataset design [12, 13] that fails to benchmark the generalization ability for the models. This overfitting issue would limit the practical applicability of these models, as they fail to adapt to the rapidly evolving landscape of deepfake techniques. Second, existing approaches lack the ability to identify the modality source of a detected deepfake. This limitation arises because these systems are trained and tested using only the final audio-visual labels, without incorporating meta-information about the individual modalities. A model able to tell which modality is fake would enhance the interpretability of the model and credibility in practice.
In this work, we propose a novel framework Multi-Stream Fusion Approach with One-Class learning (MSOC) to tackle audio-visual deepfake detection, enhancing the generalization ability and interoperability. We extend the one-class learning approach, previously proposed in unimodal contexts, to the audio-visual setting. We validate the generalization ability by re-splitting the FakeAVCeleb [12] dataset and separating the unseen algorithms into the test set. We curated four test sets (RAFV, FAFV, FAFV, Unsynced) that cover all kinds of fake categories. We will make the dataset splits and model implementation publicly available upon the publication of this paper. Our contributions are summarized as:
@@
    Extending one class learning to audio visual deepfake detection.
@@
    A multi-stream framework with audio-visual (AV), audio (A), visual (V) branches
@@
    A curated dataset for evaluating performance on unseen generation methods using FakeAVCeleb.
@@ II-A Audio-Visual Deepfake Detection
At the early stage of deepfake detection research, many studies focused on uni-modal detection models that use only audio [14] or visual [15] as input. However, uni-modal models are inherently limited to a single modality and cannot completely detect emerging deepfake videos that both audio and visual can be generated. To address this problem, recent research has started to focus on developing audio-visual deepfake detection models.
Initially, many studies have focused on explicit synchronization issues between audio and visual modalities in deepfakes. Shahzad et al. [16] argue that altering either audio or visual can desynchronize speech and lip movements in videos. Additionally, researchers have investigated representation-level inconsistency due to single modality manipulations. The modality dissonance score was introduced in [17] to quantify dissimilarities between modality features. However, these methods may struggle to detect deepfakes where both audio and video are both generated in a more consistent way, such as text-to-speech followed by lip synch [12].
Several studies also develop audio-visual representations by integrating features from unimodal feature extractors and mapping them to audio-visual targets [11, 18]. However, recent studies [10, 9] claim that using only multimodal labels can misinform unimodal feature extractor data during joint training. Ensemble models also have been studied [19, 20, 21]. They combine separate models for audio, visual, and audio-visual data and leverage the strengths of each modality-specific model to enhance overall detection accuracy.
@@ II-B Video Deepfake Detection Datasets
Existing methods are typically benchmarked on datasets such as FakeAVCeleb [12] and DFDC [13]. However, these datasets are limited in their ability to benchmark generalization since the test sets often contain the same deepfake generation algorithms as the training sets. Additionally, there is a greater variety of visual deepfake generation methods compared to audio modalities. In terms of attribute labeling, the FakeAVCeleb dataset attempts to present different categories of fakes, but the FARV category includes not only fake audio but also unsynchronized audio. This makes it difficult for methods to learn fake audio cues since they are confounded with synchronization cues. Our study proposes extended datasets and new splits to address these issues.
@@ II-C One-Class Learning For Deepfake Detection
Binary classification models work well in deepfake detection when the test data share similar distributions with the training data. However, as deepfake generation techniques are rapidly developing, deepfake attacks in practice are often unseen during the training of deepfake detection models, and these binary classification models show significantly degraded performance on unseen attacks [22]. To address this issue, Zhang et al. [23] proposed the idea of one-class learning for speech deepfake detection. The idea was to use a so-called One-Class Softmax (OC-Softmax) loss to guide the neural network to learn an embedding space where bonafide speech utterances are clustered together while fake speech utterances are pushed away from this cluster during training:
@@
Since then, many works on speech anti-spoofing have adopted the idea of one-class learning [23, 24, 25]. Results show that models trained with one-class learning can effectively identify fake as deviations from the learned bonafide embedding cluster for speech. Despite these advantages, the generalizability of one-class learning for audio-visual deepfake detection has not been thoroughly studied due to dataset limitations. This study addresses this gap by re-splitting the FakeAvCeleb dataset [12] and analyzing the effectiveness of one-class learning in audio-visual deepfake detection.
@@ III Method
We propose a Multi-Stream Fusion Approach with One-Class learning (MSOC) for audio-visual deepfake detection. This architecture consists of the audio, visual, and audio-visual branches, which are independently trained using labels specific to their respective modalities. The training of these branches also leverages the OC-Softmax loss to improve their generalization ability to unseen deepfake generation methods. During inference, score fusion is employed to integrate the decisions made by the three branches to arrive at the final classification decision. In this section, we describe the architecture, training and inference in detail.
@@ III-A Audio Model
As displayed in the green part of Fig. 1, the audio branch includes an audio feature extractor trained using an OC-Softmax loss with ground-truth labels specific to the audio modality only. The audio branch compacts real data representations of audio modality and spreads fake ones in the embedding space.
We utilize ResNet [26] as our audio feature extractor. The model processes 13-d Mel-Frequency Cepstral Coefficients (MFCC) vectors at a frame rate of 100 frames per second, which is 4 times of the visual frame rate. The audio feature extractor then produces the audio embeddings with a dimensionality of 128.
The audio model is trained with $\mathcal{L}_{AOC}^{audio}$ which is the OC-Softmax losses computed with audio features from the audio branches using audio labels.
@@ III-B Visual Model
As shown in the blue part of Fig. 1, the visual branch consists of a visual feature extractor trained with an OC-Softmax loss taking ground-truth labels regarding to the visual modality only. The visual branch tries to learn a visual embedding space where real data features are clustered while fake data features are pushed away from the cluster.
@@
ResNet. The ResNet-based visual feature extractor consists of a 3D convolutional layer, ResNet blocks, and temporal convolutional block. It captures the features of frames.
@@ III-C Audio-Visual Model
As displayed in the purple part of Fig.1, The audio-visual branch comprises OC-Softmax attached visual, audio extractors, and 3 layers of feedforward neural network. It is trained with both OC loss and cross-entropy loss. This branch focuses on compacting real data representations on each feature extractor and separating real and fake data representations across both modalities.
@@
where $\mathcal{L}_{AOC}^{av}$ and $\mathcal{L}_{VOC}^{av}$ are the OC-Softmax losses computed using audio and visual features from the audio-visual model with their respective labels. $\mathcal{L}_{CE}^{av}$ is the cross-entropy loss applied to the combined audio-visual features after the classifier, using the corresponding audio-visual labels in the audio-visual branch.
@@ IV-A Dataset
IV-A1 Dataset Split
Usually, researchers split FakeAVCeleb datasets by subject ID [9] or just randomly[20, 10]. However, these splits have limitations in testing the model’s generalizability to unseen deepfake generation methods. Thus, we split the dataset based on generation methods to evaluate model performance on unseen generation methods. While creating training, validation, and test sets, we ensured that the methods used to generate the test sets were excluded from the training and validation sets.
Training Set contains 350 videos each from categories RARV(Real), FARV, RAFV, and FAFV (excluding faceswap and faceswap-wav2lip). Validation Set contains 50 videos each from categories Real, FARV, RAFV, and FAFV (excluding faceswap and faceswap-wav2lip). For Test Set, we sampled 100 face swap (RAFV) and face swap-wav2lip (FVFA) videos not included in the training and validation sets. We generated 100 audio-only fake videos using a voice conversion library, named category FARV, due to FakeAVCeleb’s limited methods for audio fakes. It is important to note that our newly created FARV dataset is synchronized, whereas the FARV from the FakeAVCeleb[12] dataset is unsynchronized. Therefore, our FARV dataset has only one cue to detect a fake, making the unseen generation method more distinct. We also created a category Unsynced with 100 unsynchronized videos by shifting audio. Each of these four test datasets —RAFV, FVFA, FAFV, and Unsynced— consists of 100 real videos(RAFV) and 100 unseen fake videos.
@@ IV-C Comparison Methods
We adopt some existing methods from the literature for comparison. The multimodal-dissonance model [17] utilizes a modality dissonance score to detect dissimilarities between modalities. AVDF [29] concatenates audio and visual features and maps them to audio-visual labels. The multilabel method [30] is trained using both audio and visual labels to address the issue that audio-visual labels may confuse the uni-modal feature extractor. MRDF_CE and MRDF_margin [9] utilize the cross and within modality regularization.
@@ IV-D Training Details
Also, for the comparison models, we trained and tested the models on our setup from scratch. Specifically, for the multimodal-dissonance model [17], we trained for 100 epochs.
V Results
V-A Comparison with State-of-the-Art Methods
To demonstrate the effectiveness of the proposed MSOC model on unseen attacks, we compared it with other state-of-the-art audio-visual deepfake detection models. The comparison of models’ performance on test datasets is presented in table I. We can observe that state-of-the-art models perform poorly on unseen generation methods, which shows the lack of generalization ability of current state-of-the-art models. Our proposed model MSOC outperforms other models on FAFV, RAFV, and FARV test sets, with an average improvement of 7.31% in detection accuracy compared to the multilabel approach [30] over four test sets. This indicates that multi-stream architecture with OC-Softmax successfully separated bonafide and generated data by compacting embedding of bonafide data, which resulted in better generalizability than other models in all combinations of fake modality. However, it does not perform well in finding unsynchronized video. It is an expected result since our proposed model does not have an explicit module that learns audio-visual consistency.
Additionally, We compare the MSOC framework with AVOC models. Table II displays that MSOC models generally perform better than AVOC. This suggests the strength of an audio and visual branch that is only dedicated to separating real and fake in each modality.
@@ V-A Comparison with State-of-the-Art Methods
TABLE II: The table compares AVOC and MSOC models. Average % accuracy and standard deviation of four runs on each test set. The multilabel model [30] is used as a baseline.
@@
MSOC(our) Scnet-STIL $60.25\pm 2.19$ 89.88 $\pm$ 3.15 74.38 $\pm$ 5.41 45.25 $\pm$ 1.64
@@
V-B Performance Analysis of Different Branches of MSOC
Fig. 2 visualizes the distribution of scores for each branch on four categories of fake videos. The audio and visual score, OC scores $\in[-1,1]$, are calculated based on the cosine similarity between the bonafide embedding and the feature embedding of the respective modality. The audio-visual score represents the softmax probability of a video being real $\in[0,1]$, computed with audio-visual features and audio-visual labels. The figure shows that the audio-visual branch performs well when both modalities are fake (FAFV), predicting a probability close to 0 for all fake samples. However, the audio-visual branch exhibits greater confusion when only one modality is fake. The audio branch excels at distinguishing audio fake(Fig. 2(b)) and real samples(Fig. 2(a)). Also, the visual branch exhibits great performance in identifying real samples (Fig. 2(b)), although it fails to detect some fake samples (Fig. 2(a)). This highlights the benefit of leveraging audio and visual branches to enhance performance.
@@ V-C Impact of One-Class Learning
TABLE III: Comparison of models trained with and without OC softmax using different feature extractors. Average % accuracy and standard deviation of four runs. The multilabel model [30] is used as a baseline.
@@
Figure 3: T-sne visualization of concatenated audio-visual feature. The X in the figure represents the center of each category data. Better viewed in color.

### v2

@@ Abstract
This paper addresses the challenge of developing a robust audio-visual deepfake detection model. In practical use cases, new generation algorithms are continually emerging, and these algorithms are not encountered during the development of detection methods. This calls for the generalization ability of the method. Additionally, to ensure the credibility of detection methods, it is beneficial for the model to interpret which cues from the video indicate it is fake. Motivated by these considerations, we then propose a multi-stream fusion approach with one-class learning as a representation-level regularization technique. We study the generalization problem of audio-visual deepfake detection by creating a new benchmark by extending and re-splitting the existing FakeAVCeleb dataset. The benchmark contains four categories of fake videos (Real Audio-Fake Visual, Fake Audio-Fake Visual, Fake Audio-Real Visual, and Unsynchronized videos). The experimental results demonstrate that our approach surpasses the previous models by a large margin. Furthermore, our proposed framework offers interpretability, indicating which modality the model identifies as more likely to be fake. The source code is released at https://github.com/bok-bok/MSOC.
@@ I Introduction
Recent research focuses mainly on the fusion of features of both modalities to improve the detection performance on audio-visual deepfake datasets [9, 10, 11]. By leveraging the complementary nature of audio and visual data, these approaches effectively improve their accuracy in identifying manipulated content. However, two issues are not well explored: First, the existing deep learning models may overfit to the specific fake generation methods present in the training data, leading to poor generalization when confronted with unseen deepfake generation algorithms in real-world scenarios. This could be attributed to the existing dataset design [12, 13] that does not benchmark the generalization ability for the models. This overfitting issue would limit the practical applicability of these models, as they fail to adapt to the rapidly evolving landscape of deepfake techniques. Second, existing approaches lack the ability to identify the modality source of a detected deepfake. This limitation arises because these systems are trained and tested using only the final audio-visual labels, without incorporating meta-information about the individual modalities. A model able to tell which modality is fake would enhance the interpretability and credibility in practice.
In this work, we propose a novel framework Multi-Stream Fusion Approach with One-Class Learning (MSOC) to tackle audio-visual deepfake detection, enhancing the generalization ability and interoperability. We extend the one-class learning approach, previously proposed in uni-modal contexts, to the audio-visual setting. We validate the generalization ability by resplitting the FakeAVCeleb [12] dataset and separating the unseen algorithms into the test set. We curated four test sets (RAFV, FAFV, FAFV, Unsynced) that cover all kinds of fake categories. We will make the dataset splits and model implementation publicly available upon the publication of this paper. Our contributions are summarized as:
@@
    Extending one class learning from uni-modal to audio-visual deepfake detection;
@@
    A multi-stream framework with audio-visual (AV), audio (A), and visual (V) branches;
@@
    A curated dataset for evaluating performance on unseen generation methods based on FakeAVCeleb.
@@ II-A Audio-Visual Deepfake Detection
At the early stage of deepfake detection research, many studies focused on uni-modal detection models that use only audio [14, 15] or visual [16] as input. However, uni-modal models are inherently limited to a single modality and cannot completely detect emerging deepfake videos that both audio and visual can be generated. To address this problem, recent research has started to focus on developing audio-visual deepfake detection models.
Initially, many studies have focused on explicit synchronization issues between audio and visual modalities in deepfakes. Shahzad et al. [17] argue that altering either audio or visual can desynchronize speech and lip movements in videos. In addition, researchers have investigated the representation-level inconsistency due to single-modality manipulations. The modality dissonance score was introduced in [18] to quantify the dissimilarities between the modality features. However, these methods may struggle to detect deepfakes where both audio and video are both generated in a more consistent way, such as text-to-speech followed by lip synch [12].
Several studies also develop audio-visual representations by integrating features from uni-modal feature extractors and mapping them to audio-visual targets [11, 19]. However, recent studies [10, 9] claim that using only multimodal labels can misinform the data from the unimodal feature extractor during joint training. Ensemble models have also been studied [20, 21, 22]. They combine models for audio, visual, and audio-visual data and leverage the strengths of each modality-specific model to enhance overall detection accuracy.
@@ II-B Video Deepfake Detection Datasets
Existing methods are typically benchmarked on datasets such as FakeAVCeleb [12] and DFDC [13]. However, these datasets are limited in their ability to benchmark generalization since the test sets often contain the same deepfake generation algorithms as the training sets. Additionally, there is a greater variety of visual deepfake generation methods compared to audio modalities. In terms of attribute labeling, the FakeAVCeleb dataset attempts to present different categories of fakes, but the FARV category includes not only fake audio but also unsynchronized audio. This makes it difficult for methods to learn fake audio cues, since they are confounded with synchronization cues. Our study proposes extended datasets and new partitions to address these issues.
@@ II-C One-Class Learning For Deepfake Detection
Binary classification models work well in deepfake detection when the test data share similar distributions with the training data. However, since deepfake generation techniques are rapidly developing, deepfake attacks in practice are often unseen during training of deepfake detection models, and these binary classification models show significantly degraded performance on unseen attacks [23]. To address this issue, Zhang et al. [14] proposed the idea of one-class learning for speech deepfake detection. The idea was to use a so-called One-Class Softmax (OC-Softmax) loss to guide the neural network to learn an embedding space where bonafide speech utterances are clustered together while fake speech utterances are pushed away from this cluster during training:
@@
Since then, many works on speech anti-spoofing have adopted the idea of one-class learning [24, 15, 25]. The results show that models trained with one-class learning can effectively identify fakes as deviations from the learned bonafide embedding cluster for speech. Despite these advantages, the generalizability of one-class learning for audio-visual deepfake detection has not been thoroughly studied due to dataset limitations. This study addresses this gap by re-splitting the FakeAVCeleb dataset [12] and analyzing the effectiveness of one-class learning in audio-visual deepfake detection.
@@ III Method
We propose a Multi-Stream Fusion Approach with One-Class learning (MSOC) for audio-visual deepfake detection. This architecture consists of the audio, visual, and audio-visual branches, which are independently trained using labels specific to their respective modalities. The training of these branches also leverages the OC-Softmax loss to improve their generalization ability to unseen deepfake generation methods. During inference, score fusion is used to integrate the decisions made by the three branches to arrive at the final classification decision. In this section, we describe architecture, training, and inference in detail.
@@ III-A Audio Model
As displayed in the green part of Fig. 1, the audio branch includes an audio feature extractor trained using an OC-Softmax loss with ground truth labels specific only to the audio modality. The audio branch compacts real data representations of audio modality and spreads fake ones in the embedding space.
We utilize ResNet [26] as our audio feature extractor. The model processes 13-d Mel-Frequency Cepstral Coefficients (MFCC) vectors at a frame rate of 100 frames per second, which is 4 times the visual frame rate. The audio feature extractor then produces the audio embeddings with a dimensionality of 128.
The audio model is trained with $\mathcal{L}_{AOC}^{audio}$ which is the OC-Softmax losses computed with audio features of the audio branches using audio labels.
@@ III-B Visual Model
As shown in the blue part of Fig. 1, the visual branch consists of a visual feature extractor trained with an OC-Softmax loss, taking ground-truth labels regarding the visual modality only. The visual branch tries to learn a visual embedding space where real data features are clustered while fake data features are pushed away from the cluster.
@@
ResNet. The ResNet-based visual feature extractor consists of a 3D convolutional layer, ResNet blocks, and a temporal convolutional block. It captures the features of frames.
@@ III-C Audio-Visual Model
As shown in the purple part of Fig. 1, the audio-visual branch consists of OC-Softmax integrated with visual and audio extractors, followed by three layers of a feedforward neural network. It is trained with both OC loss and cross-entropy loss. This branch focuses on compacting real-data representations on each feature extractor and separating real- and fake-data representations across both modalities.
@@
where $\mathcal{L}_{AOC}^{av}$ and $\mathcal{L}_{VOC}^{av}$ are the OC-Softmax losses computed using audio and visual features from the audio-visual model with their respective labels. $\mathcal{L}_{CE}^{av}$ is the cross-entropy loss applied to the combined audio-visual features after the classifier, using the corresponding labels in the audio-visual branch.
@@ IV-A Dataset
Previous works typically split the FakeAVCeleb dataset by subject ID [9] or randomly [21, 10]. However, these splits have limitations in assessing the model’s generalizability to unseen deepfake generation methods. In this paper, we propose a new split mechanism: we split the dataset based on generation methods to evaluate the performance on unseen methods. During the creation of training, validation, and test sets, we ensured that the generation methods used in the test sets were excluded from the training and validation sets.
Our Training set contains 350 videos each from categories RARV (Real), FARV, RAFV, and FAFV (excluding faceswap and faceswap-wav2lip). Validation set contains 50 videos each from categories Real, FARV, RAFV, and FAFV (excluding faceswap and faceswap-wav2lip). For Test set, we sampled 100 face swap (RAFV) and face swap-wav2lip (FVFA) videos not included in the training and validation sets. We generated 100 audio-only fake videos using a voice conversion library, named category FARV, due to FakeAVCeleb’s limited methods for audio fakes. It is important to note that our newly created FARV dataset is synchronized, whereas the FARV from the FakeAVCeleb[12] dataset is unsynchronized. Therefore, our FARV dataset has only one cue to detect a fake, making the unseen generation method more distinct. We also created a Unsynced category with 100 unsynchronized videos by shifting audio. Each of these four test datasets — RAFV, FVFA, FAFV, and Unsynced— consists of 100 real videos (RARV) and 100 unseen fake videos.
@@ IV-C Comparison Methods
We adopt some existing methods from the literature for comparison. The multimodal-dissonance model [18] utilizes a modality dissonance score (a distance between audio and visual features) to detect dissimilarities between modalities. AVDF [29] simply concatenates audio and visual features and maps them directly to audio-visual labels. The multilabel method [30] is trained using both audio and visual labels to address the issue that audio-visual labels may confuse the uni-modal feature extractors. MRDF_CE and MRDF_margin [9] utilize the cross- and within-modality regularization to maintain the unique features and differences of each modality during multimodal representation learning.
@@ IV-D Training Details
Also, for the comparison models, we trained and tested the models in our set-up from scratch for a fair comparison. Specifically, for the multimodal-dissonance model [18], we trained for 100 epochs.
@@
TABLE II: The table compares AVOC and MSOC models. Average accuracy (%) and standard deviation of four runs on each test set. The multilabel model [30] is used as a baseline.
@@
MSOC (Ours) SCNet-STIL $60.25\pm 2.19$ 89.88 $\pm$ 3.15 74.38 $\pm$ 5.41 45.25 $\pm$ 1.64
@@
V Results
V-A Comparison with State-of-the-Art Methods
To demonstrate the effectiveness of the proposed MSOC model on unseen attacks, we compared it with other state-of-the-art audio-visual deepfake detection models. The comparison of models’ performance on test datasets is presented in Table I. We can observe that state-of-the-art models perform poorly on unseen generation methods, which shows their lack of generalization ability. Our proposed model MSOC outperforms other models on FAFV, RAFV, and FARV test sets. This indicates that multi-stream architecture with OC-Softmax successfully separated bonafide and generated data by compacting embedding of bonafide data, which resulted in better generalizability than other models in all combinations of fake modality.
As shown in the last column of Table I, all models perform poorly (close to random guessing) when identifying unsynchronized videos, which should be clearly recognized as fake. This is the first time these models have been tested on this unsynchronization benchmark, and our model exhibits general characteristics similar to existing fusion-based methods. The results suggest that training the audio and visual encoders with real/fake labels alone is insufficient to capture synchronization. We believe that incorporating an explicit module to learn audio-visual synchronization [31, 32] could address this issue, but we leave this for future work.
Additionally, we compare the MSOC framework with AVOC models. Table II shows that MSOC models generally perform better than AVOC. This suggests the strength of an audio and visual branch that is only dedicated to separating real and fake in each modality.
@@ V-A Comparison with State-of-the-Art Methods
V-B Performance Analysis of Different Branches of MSOC
Fig. 2 visualizes the distribution of scores for each branch on four categories of fake videos. The audio and visual score, OC scores $\in[-1,1]$, are calculated based on the cosine similarity between the bonafide embedding and the feature embedding of the respective modality. The audio-visual score represents the softmax probability that a video is real $\in[0,1]$, calculated with audio-visual characteristics and audio-visual labels. The figure shows that the audio-visual branch performs well when both modalities are fake (FAFV), predicting a probability close to 0 for all fake samples. However, the audio-visual branch exhibits greater confusion when only one modality is fake. The audio branch excels at distinguishing audio fake(Fig. 2(b)) and real samples(Fig. 2(a)). Also, the visual branch exhibits great performance in identifying real samples (Fig. 2(b)), although it fails to detect some fake samples (Fig. 2(a)). This highlights the benefit of using both the audio and visual branches.
@@ V-C Impact of One-Class Learning
TABLE III: Comparison of models trained with and without OC softmax using different feature extractors. Average accuracy (%) and standard deviation of four runs. The multilabel model [30] is used as a baseline.
@@
Figure 3: t-SNE visualization of concatenated audio-visual feature. The cross “X” in the figure represents the center of the data for each category. Better viewed in color.